Find values of a and k for which each parabola contains the given points
y-k = a(x+3)^2 ; (-5,1) (1,7)

Answer by lwsshak3(11628)

Find values of a and k for which each parabola contains the given points
y-k = a(x+3)^2 ; (-5,1) (1,7)
**
y-k = a(x+3)^2
y=a(x+3)^2+k
set up two equations with coordinates of given points:
1=a(-5+3)^2+k
7=a(1+3)^2+k
..
1=a(-2)^2+k
7=a(4)^2+k
..
1=4a+k
7=16a+k
..
4=16a+4k
7=16a+k
..
-3=3k
k=-1
..
1=4a-1
4a=2
a=1/2
check:(using point (1,7)
7-(-1)=1/2(1+3)^2
8=8